A sonic depth gauge is placed 5.8 m above the ground. An ultrasound pulse sent downward reflects off snow and reaches the device 0.031 seconds after it was emitted. The air temperature is −20°C.

(a)
How far (in m) is the surface of the snow from the device? (Round your answer to at least three significant figures.)


(b)
How deep (in m) is the snow?
1
Expert's answer
2019-04-30T09:58:11-0400

The speed of air at −20°C is equal to 318.8 m/s

(a) Distance to snow is equal to


"\frac {318.8 m/s \times 0.031 s} {2}=4.941 m"

(b) How deep (in m) is the snow?

5.8 m-4.941 m=0.859 m


Answer:

(a) 4.941 m

(b) 0.859 m